# Multi-Excitation Projective Simulation with a Many-Body Physics Inspired Inductive Bias

### v2

@@ IV.2 Formalization
Consider a simple 2-layer setting, with 4 clips in each layer, see Figure 3: $V=L_{1}\cup L_{2}$, with $L_{1}=\{c_{1},c_{2},c_{3},c_{4}\}$ and $L_{2}=\{c^{\prime}_{1},c^{\prime}_{2},c^{\prime}_{3},c^{\prime}_{4}\}$. We only consider h-values with the same number of incoming and outgoing excitation numbers, and let no more than two excitations interact. That means $IO=\{(1,1),(2,2)\}$. Our current excitation configuration is $\{c_{1},c_{2},c_{3}\}$, meaning that we currently have an excitation in each of the clips $c_{1}$, $c_{2}$, and $c_{3}$.
@@
This list gets turned into probabilities, in our example by applying the softmax-function to the full list. Say, we sample $h_{(2,2)}(\{c_{2},c_{3}\},\{c_{1},c^{\prime}_{1}\})$ and apply it to our current configuration $\{c_{1},c_{2},c_{3}\}$. First, we remove the excitations in $c_{2}$ and $c_{3}$, giving us the configuration $\{c_{1}\}$. Next, we put excitations into $c_{1}$ and $c^{\prime}_{1}$. However, $c_{1}$ already carries an excitation. We just keep this excitation as it is. So our next excitation configuration is $\{c_{1},c^{\prime}_{1}\}$. Note that our rule for dealing with already occupied clips led to a reduction in the total number of excitations.
Our layered Inductive Biases 2FF and 2SF differ from the previous situation in that the relevant many-body h-values are only items 1 and 4 from the numbered list above. Now, say that we sampled $h_{(2,2)}(\{c_{1},c_{2}\},\{c^{\prime}_{2},c^{\prime}_{3}\})$ and apply it to our current configuration $\{c_{1},c_{2},c_{3}\}$. First, we remove the excitations in $c_{1}$ and $c_{2}$, giving us the configuration $\{c_{3}\}$. Next, we insert excitations in $c^{\prime}_{2},c^{\prime}_{3}$, giving us the full next excitation configuration $\{c^{\prime}_{2},c^{\prime}_{3},c_{3}\}$. We observe that while the feed-forward condition forces all excitations that move to move one layer forward, it allows excitations to stay behind in their old clip in the old layer. Consider now an additional layer $L_{3}$. While Inductive Bias 2FF allows us to continue with any transition $C_{\mathrm{in}}\rightarrow C_{\mathrm{out}}$ that has $C_{\mathrm{in}}\subset\{c^{\prime}_{2},c^{\prime}_{3}\}$ or $C_{\mathrm{in}}=\{c_{3}\}$, Inductive Bias 2SF forces us to remove $c_{3}$ first. For Inductive Bias 2SF, the relevant many-body h-values are therefore just those of the form $h_{(1,1)}(c_{3},c^{\prime}_{j})$ with $j=1,\dots 4$.
Now, consider Inductive Bias 2DP acting on $\{c_{1},c_{2},c_{3}\}$. While it has the same list of relevant $h$-values as Inductive Bias 2FF, it applies the transitions another way. Again, assume that we sampled $h_{(2,2)}(\{c_{1},c_{2}\},\{c^{\prime}_{2},c^{\prime}_{3}\})$ and apply it to $\{c_{1},c_{2},c_{3}\}$. Again, we remove the ingoing excitations $c_{1}$ and $c_{2}$, giving us $\{c_{3}\}$. Next, we insert excitations in $c^{\prime}_{2}$ and $c^{\prime}_{3}$, giving us $\{c^{\prime}_{2},c^{\prime}_{3},c_{3}\}$. Furthermore, $c_{3}$ is neither an ingoing nor an outgoing clip of $h_{(2,2)}(\{c_{1},c_{2}\},\{c^{\prime}_{2},c^{\prime}_{3}\})$, so we discard its excitation. This gives us the full next excitation configuration $\{c^{\prime}_{2},c^{\prime}_{3}\}$. As we see, Inductive Bias 2DP enforces that after a transition, all excitations are found in the same layer.
Figure 3: An example illustrating random walk steps under different inductive biases, compare with Example 11. Excited clips are shown in red. The sampled hyperedge is shown in blue. Subfigure a) shows a deliberation step which is only allowed under Inductive Bias 1, because its codomain is in two layers. Also, it shows that an excitation moving into an occupied clip gets discarded. Subfigure b) shows a typical transition under Inductive Biases 2FF and 2SF. Notice that it can lead to excitations being spread over several layers. Subfigure c) shows a typical transition under Inductive Bias 2DP. It discards uninvolved excitations, and therefore only contains excitations in the codomain of the hyperedge.
IV.3 Complexity Estimates
@@
From this definition, one can quickly see that unrestricted mePS agents have several costs associated to them that scale (at least) exponentially in the number of clips $|V|$.
@@
Consider a mePS agent obeying Inductive Bias 1, 2FF, 2SF, or 2DP. Define $\max I:=\max\{i\ |\ \exists o:(i,o)\in IO\}$ and $\max O:=\max\{o\ |\ \exists i:(i,o)\in IO\}$, as well as $\max IO:=\max\{i+o\ |\ (i,o)\in IO\}$. Then the number of trainable parameters is $\mathcal{O}(\max I\cdot\max O\cdot|V|^{\max IO})$.
@@
First, we note that $|IO|\leq\max I\cdot\max O$. For each $(i,o)$, let us bound the number of $C_{I},C_{O}\in\mathcal{P}(V)$ for the many-body h-values $h_{(i,o)}(C_{I},C_{O})$. Using binomial coefficients and Inductive Bias 1, this number is upper-bounded by $\begin{pmatrix}|V|\\ i\end{pmatrix}\cdot\begin{pmatrix}|V|\\ o\end{pmatrix}\leq|V|^{i}|V|^{o}=|V|^{i+o}\leq|V|^{\max IO}$. So, the number of many-body h-values for each $(i,o)$ is upper-bounded by $|V|^{\max IO}$. Since we have at most $\max I\cdot\max O$ choices for $(i,o)$, the total number of h-values is upper bounded by $\max I\cdot\max O\cdot|V|^{\max IO}$. The Inductive Biases 2 have even fewer many-body h-values than Inductive Bias 1 alone would allow. ∎
@@
So far, our resource estimates do not distinguish between our inductive biases. However, we already hinted at the fact that they have crucial complexity differences with regard to the maximal deliberation time. To see the difference, we derive upper bounds on the deliberation time.
At first, we point out that Inductive Bias 1 still allows for deliberations to be arbitrarily long: For any attainable excitation configuration $\{c_{m_{1}},\dots,c_{m_{x}}\}$, if it is possible to combine transitions relevant for this configuration to a cycle leading again to $\{c_{m_{1}},\dots,c_{m_{x}}\}$, then this cycle can lead to arbitrarily long deliberation times. A simple class of Inductive Bias 1 agents that have such cycles is presented in Proposition 20 in Appendix C. One important consequence of our feed-forward conditions is that they explicitly break such cycles.
In that regard, we first establish that our Inductive Bias 2FF indeed leads to a finite upper bound on the deliberation time:
Proposition 16.
Assume Inductive Bias 2FF for a layered mePS agent with layers $(L_{1},\dots,L_{D})$. Then the deliberation time (i.e. the total number of random walk steps) is upper-bounded by $\prod_{j=1}^{D-1}(|L_{j}|+1)$.
The proof is in Appendix C. While the upper bound in Proposition 16 is finite, it is exponential in the depth $D$. However, there exist scenarios in which there is also an exponential lower bound on the maximal deliberation time, see Proposition 22 in Appendix C. To get this exponential scaling, it is enough to consider examples with $IO=\{(1,2)\}$, despite the fact that this $IO$ only contains small $i$ and $o$.
One key point of the proofs is that for all these scenarios, we require $(i,o)$ with $o>i$ such that we can start an “avalanche” of excitations. However, if we choose our inductive bias on $IO$ such that the number of excitations cannot increase, we find a much better upper bound, which is essentially $\mathrm{width}\times\mathrm{depth}^{2}$:
Proposition 17.
Proof.
Furthermore, the proof of Proposition 16 relies on allowing excitations to be moved to the deepest layers first. Inductive Bias 2SF explicitly forces the shallowest excitations to move first instead, also resulting in a much more efficient upper bound:
Proposition 18.
Proof.
Proposition 19.
Proof.
More complexity estimates can be found in Appendix C.
@@ V.1 Invasion Game With Distraction
The different few-body inductive biases we will consider play a large role in determining the complexity of the agent, illustrating the general idea behind inductive biases: adapting the agent’s architecture and complexity to suit the task environment. For this purpose, our goal is to construct an environment that cannot be solved through consideration of only one excitation at a time, but can be solved perfectly by looking at two excitations and is “overcomplicated” when looking at three excitations simultaneously.
An environment achieving this goal is constructed as follows: Each round, the agent obtains a percept/observation of the form $(v_{1},v_{2},v_{3})\in\{0,1,2\dots,9\}^{\times 3}$. Here, each entry $v_{j}$ corresponds to a value of an observable $j$. For each observable $j$ and each value $v_{j}$, we associate a clip denoted “$\mathrm{obs_{j}:v_{j}}$” in the percept layer of our two-layer agents. Therefore, each percept is coupled in by putting three excitations into the corresponding clips in the percept layer. As an action, the agent has to pick exactly one of two doors. These two choices are represented by clips “$a=0$” and “$a=1$” in the action layer of the agent. The action gets coupled out as soon as there is an excitation on one of the action clips.
@@
For $i\in\{1,2,3\}$, we consider two-layered (percept+action layer) agents with many-body inductive bias using $IO=\{(i,1)\}$, and use percept and action layers as described above. Since we directly couple out actions, there is no difference between the different Inductive Biases 2. This allows us to directly focus on the difference caused by different choices of $i$.
@@ V.2 Deceptive Invasion Game
It is again expected that the $(2,1)$-agent will reach the optimal policy the quickest, followed by the $(3,1)$-agent taking more time due to the processing of irrelevant percepts represented by the third observable. Due to the more complex reward structure of this environment compared with the Invasion Game With Distraction, the $(1,1)$-agent’s task will become even more impossible than it already was because it has to cut through the attacker’s deception on top of the already present obstacles in the Invasion Game With Distraction environment.
@@ V.3 Computer Maintenance
Symptoms
Components
Causes
Fixes
PC overheating: 1
CPU: 13
physical damage: 18
replace components: 9
files disappearing: 2
SSD: 14
software damage: 19
install missing software: 10
visible markings on components: 3
MoBo: 15
malware: 20
cooldown computer: 11
unexpected shutdowns: 4
PSU: 16
faulty: 21
run antivirus: 12
slow performance: 5
OS: 17
not connected: 22
@@
The agent must learn to navigate and solve 11 different scenarios in this work, which we constrained to have at most 2 elements per category to better demonstrate the savings accrued from reducing the size of the ECM. An example scenario is detailed in Figure 7. It is immediately clear from this example scenario that the feedback loop between the MoBo and the SSD demands that these components and the associated causes all be treated together if the agent hopes to solve the problem – something the mePS agent is better suited to handle. The only hope that the single-excitation agent has of solving the scenario is to exponentially inflate the size of their ECM so that all possible combinations of the environment variables are distributed in all clips. If each environment variable can take sufficiently many values, the previous strategy will undoubtedly fail since the random walk path in the agent’s ECM that represents the correct solution to a given scenario will have a vanishingly small probability of occurring.
@@
An inductive bias agent will be represented with the following notation: $\{n_{s},[n_{hc},n_{c}],[n_{ac},n_{f}]\}$, which we call an inductive bias configuration. Here, $n_{s}$ is the many-body cutoff on the number of symptoms in the inductive bias agent’s percept layer; $n_{hc}$, the cutoff on the number of components in the hidden layer; $n_{c}$, the cutoff on the number of causes in the hidden layer; $n_{ac}$, the cutoff on the number of components in the action layer; and $n_{f}$, the cutoff on the number of fixes in the action layer. Since we are discarding leftover excitations, and because we have a layered, feed-forward architecture, the inductive bias used here corresponds with Inductive Bias 2DP. We can calculate the number of learning parameters $N_{l}$ for each agent from Equation (10) in Appendix A, which we will use throughout the rest of this section.
@@
All agents for each of their layers use the softmax function with $\beta=1/2$ to convert $h$-values to probabilities and all $h$-values are initialized to 1. The results of the training are shown in Figure 8. We can see that the inductive bias agent achieves near-optimal values for both the hypothesis and plausibility rewards using far fewer steps but requiring more episodes than the unrestricted agent does, as seen in the total step number per episode curve, which better illustrates the savings from the inductive bias as the two average reward plots do not show how many total steps have elapsed during training. The total step number per episode curve also shows roughly exponential decay, which is indicative of the agent first trying out scenario configurations randomly and then always picking the right configuration immediately. To quantify the difference in step number to the optimal reward, we calculated the total average steps taken over 200 episodes for both agents: the inductive bias agent takes roughly 5474 steps while the unrestricted agent takes roughly 8168 steps. Together with the complexity-theoretic results from Section IV.3, we only expect this difference in step number to grow. Although the unrestricted agent does converge to the optimal rewards in fewer episodes, it clearly takes more steps to do so because it has roughly 16 times the number of trainable parameters as the inductive bias agent has. The fact that we get comparable performance from the inductive bias agent using far fewer parameters justifies the relatively small fluctuations around the optimal value for each of the reward curves, which are expected to decrease as more agents are included in the average.
@@
VI First Steps Towards Quantum mePS

### v3

@@ IV A Physics-Inspired Inductive Bias
Consider a simple 2-layer setting, with 4 atomic clips in each layer, see Figure 3: $V=L_{1}\cup L_{2}$, with $L_{1}=\{c_{1},c_{2},c_{3},c_{4}\}$ and $L_{2}=\{c^{\prime}_{1},c^{\prime}_{2},c^{\prime}_{3},c^{\prime}_{4}\}$. We only consider h-values with the same number of incoming and outgoing excitation numbers, and let no more than two excitations interact. That means $IO=\{(1,1),(2,2)\}$. Our current excitation configuration is $\{c_{1},c_{2},c_{3}\}$, meaning that we currently have an excitation in each of the atomic clips $c_{1}$, $c_{2}$, and $c_{3}$.
@@
This list gets turned into probabilities, in our example by applying the softmax-function to the full list. Say, we sample $h_{(2,2)}(\{c_{2},c_{3}\},\{c_{1},c^{\prime}_{1}\})$ and apply it to our current configuration $\{c_{1},c_{2},c_{3}\}$. First, we remove the excitations in $c_{2}$ and $c_{3}$, giving us the configuration $\{c_{1}\}$. Next, we put excitations into $c_{1}$ and $c^{\prime}_{1}$. However, $c_{1}$ already carries an excitation. We just keep this excitation as it is. So our next excitation configuration is $\{c_{1},c^{\prime}_{1}\}$. Note that our rule for dealing with already occupied atomic clips led to a reduction in the total number of excitations.
Our layered Inductive Bias 2FF differs from the previous situation in that the relevant many-body h-values are only items 1 and 4 from the numbered list above. Now, say that we sampled $h_{(2,2)}(\{c_{1},c_{2}\},\{c^{\prime}_{2},c^{\prime}_{3}\})$ and apply it to our current configuration $\{c_{1},c_{2},c_{3}\}$. First, we remove the excitations in $c_{1}$ and $c_{2}$, giving us the configuration $\{c_{3}\}$. Next, we insert excitations in $c^{\prime}_{2},c^{\prime}_{3}$, giving us the full next excitation configuration $\{c^{\prime}_{2},c^{\prime}_{3},c_{3}\}$. We observe that while the feed-forward condition forces all excitations that move to move one layer forward, it allows excitations to stay behind in their old atomic clip in the old layer. Consider now an additional layer $L_{3}$. Inductive Bias 2FF allows us to continue with any transition $C_{\mathrm{in}}\rightarrow C_{\mathrm{out}}$ that has $C_{\mathrm{in}}\subset\{c^{\prime}_{2},c^{\prime}_{3}\}$ or $C_{\mathrm{in}}=\{c_{3}\}$.
Physically, this layered structure corresponds to situations encountered, for example, in integrated photonics chips performing quantum computation with several photons: The photons move forward in the lateral direction, but perform a quantum walk in the transversal direction [46, 47]. A common noise source of such chips is that photons get absorbed by the environment.
@@
From this definition, one can quickly see that unrestricted mePS agents have several costs associated to them that scale (at least) exponentially in the number of atomic clips $|V|$.
@@
Consider a mePS agent obeying Inductive Bias 1 or 2FF. Define $\max I:=\max\{i\ |\ \exists o:(i,o)\in IO\}$ and $\max O:=\max\{o\ |\ \exists i:(i,o)\in IO\}$, as well as $\max IO:=\max\{i+o\ |\ (i,o)\in IO\}$. Then the number of trainable parameters is $\mathcal{O}(\max I\cdot\max O\cdot|V|^{\max IO})$.
@@
First, we note that $|IO|\leq\max I\cdot\max O$. For each $(i,o)$, let us bound the number of $C_{I},C_{O}\in\mathcal{P}(V)$ for the many-body h-values $h_{(i,o)}(C_{I},C_{O})$. Using binomial coefficients and Inductive Bias 1, this number is upper-bounded by $\begin{pmatrix}|V|\\ i\end{pmatrix}\cdot\begin{pmatrix}|V|\\ o\end{pmatrix}\leq|V|^{i}|V|^{o}=|V|^{i+o}\leq|V|^{\max IO}$. So, the number of many-body h-values for each $(i,o)$ is upper-bounded by $|V|^{\max IO}$. Since we have at most $\max I\cdot\max O$ choices for $(i,o)$, the total number of h-values is upper bounded by $\max I\cdot\max O\cdot|V|^{\max IO}$. Inductive Bias 2FF has even fewer many-body h-values than Inductive Bias 1 alone would allow. ∎
@@
While Inductive Bias 2FF is naturally adapted to layered ECMs and only has an amount of trainable parameters polynomial in the number of atomic clips, it still allows for some pathological many-body hypergraphs for which there exist random walk paths that have a length exponential in the number of layers. Specifically, in Proposition 19 in Appendix D we consider many-body hypergraphs with $o>i=1$. These can start an “avalanche” of new excitations every time one tries to move one excitation forward. If one removes these excitations in the worst order (from deep to shallow), we prove that these random walk paths have an exponential length.
While we do not expect these worst-case avalanche paths to matter in practice (an agent which indicates that almost all possible concepts matter is not very helpful), we still propose modifications of Inductive Bias 2FF which explicitly make such paths impossible.
Specifically, in Appendix B, we introduce Inductive Bias 2SF and 2DP. Inductive Bias 2SF differs from 2FF by demanding that the shallowest excitations get removed first. In Appendix D, we prove that Inductive Bias 2SF has a maximal random walk length of $\mathcal{O}(D\cdot\max_{j}{|L_{j}|})$, which is linear in both the width and depth of the ECM.
Such agents can be interpreted as forgetting the oldest facts/concepts/atomic clips first. If one wishes such old atomic clips (or rather, their semantics) to persist, one needs to model $E^{(i,o)}$ such that it allows one to copy-paste these atomic clips into deeper layers.
Inductive Bias 2DP is the harshest one, it discards all passive excitations which did not contribute to a sampled transition. This is an agent which only keeps in mind atomic clips which are immediately relevant, but as a tradeoff its random walk lengths are upper-bounded by $\mathcal{O}(D)$, as we also prove in Appendix D.
However, we consider exponential excitation avalanches to be pathological as they are not helpful for interpretability, and therefore restrict our attention in the main text to Inductive Bias 2FF, which is the most flexible.
@@ V.1 Invasion Game With Distraction
The different few-body inductive biases we will consider play a large role in determining the complexity of the agent, illustrating the general idea behind inductive biases: adapting the agent’s biases and complexity to suit the task environment. For this purpose, our goal is to construct an environment that cannot be solved through consideration of only one excitation at a time, but can be solved perfectly by looking at two excitations and is “overcomplicated” when looking at three excitations simultaneously.
An environment achieving this goal is constructed as follows: Each round, the agent obtains a percept/observation of the form $(v_{1},v_{2},v_{3})\in\{0,1,2\dots,9\}^{\times 3}$. Here, each entry $v_{j}$ corresponds to a value of an observable $j$. For each observable $j$ and each value $v_{j}$, we associate an atomic clip denoted “$\mathrm{obs_{j}:v_{j}}$” in the percept layer of our two-layer agents. Therefore, each percept is coupled in by putting three excitations into the corresponding atomic clips in the percept layer. As an action, the agent has to pick exactly one of two doors. These two choices are represented by atomic clips “$a=0$” and “$a=1$” in the action layer of the agent. The action gets coupled out as soon as there is an excitation on one of the atomic action clips.
@@
For $i\in\{1,2,3\}$, we consider two-layered (percept+action layer) agents with many-body inductive bias using $IO=\{(i,1)\}$, and use percept and action layers as described above. This allows us to directly focus on the difference caused by different choices of $i$.
@@ V.2 Deceptive Invasion Game
It is again expected that the $(2,1)$-agent will reach the optimal policy the quickest, followed by the $(3,1)$-agent taking more time due to the processing of irrelevant percepts represented by the third observable. Due to the more complex reward structure of this environment compared with the Invasion Game With Distraction, the $(1,1)$-agent’s task will become even less feasible than it already was because it has to cut through the attacker’s deception on top of the already present obstacles in the Invasion Game With Distraction environment.
@@ V.3 Computer Maintenance
Symptoms Components Causes Fixes
PC overheating: 1 CPU: 13 physical damage: 18 replace components: 9
files disappearing: 2 SSD: 14 software damage: 19 install missing software: 10
visible markings on components: 3 MoBo: 15 malware: 20 cooldown computer: 11
unexpected shutdowns: 4 PSU: 16 faulty: 21 run antivirus: 12
slow performance: 5 OS: 17 not connected: 22
@@
The agent must learn to navigate and solve 11 different scenarios in this work, which we constrained to have at most 2 elements per category to better demonstrate the savings accrued from reducing the size of the ECM. An example scenario is detailed in Figure 7. It is immediately clear from this example scenario that the feedback loop between the MoBo and the SSD demands that these components and the associated causes all be treated together if the agent hopes to solve the problem – something the mePS agent is better suited to handle. The only hope that the single-excitation agent has of solving the scenario is to exponentially inflate the size of their ECM so that all possible combinations of the environment variables are distributed in all atomic clips. If each environment variable can take sufficiently many values, the previous strategy will undoubtedly fail since the random walk path in the agent’s ECM that represents the correct solution to a given scenario will have a vanishingly small probability of occurring.
@@
An inductive bias agent will be represented with the following notation: $\{n_{s},[n_{hc},n_{c}],[n_{ac},n_{f}]\}$, which we call an inductive bias configuration. Here, $n_{s}$ is the many-body cutoff on the number of symptoms in the inductive bias agent’s percept layer; $n_{hc}$, the cutoff on the number of components in the hidden layer; $n_{c}$, the cutoff on the number of causes in the hidden layer; $n_{ac}$, the cutoff on the number of components in the action layer; and $n_{f}$, the cutoff on the number of fixes in the action layer. Since we are discarding leftover excitations, and because we have a layered, feed-forward architecture, the inductive bias used here corresponds to Inductive Bias 2DP from Appendix B. It differs from Inductive Bias 2FF only in the fact that it discards passive excitations, i.e. $C_{\mathrm{out}}:=C^{(o)}_{\mathrm{out}}$. We can calculate the number of learning parameters $N_{l}$ for each agent from Equation (10) in Appendix A, which we will use throughout the rest of this section.
@@
All agents for each of their layers use the softmax function with $\beta=1/2$ to convert $h$-values to probabilities and all $h$-values are initialized to 1. The results of the training are shown in Figure 8. We can see that the inductive bias agent achieves near-optimal values for both the hypothesis and plausibility rewards using far fewer steps but requiring more episodes than the unrestricted agent does, as seen in the total step number per episode curve, which better illustrates the savings from the inductive bias as the two average reward plots do not show how many total steps have elapsed during training. The total step number per episode curve also shows roughly exponential decay, which is indicative of the agent first trying out scenario configurations randomly and then always picking the right configuration immediately. To quantify the difference in step number to the optimal reward, we calculated the total average steps taken over 200 episodes for both agents: the inductive bias agent takes roughly 5474 steps while the unrestricted agent takes roughly 8168 steps. Together with the complexity-theoretic results from Appendix D, we only expect this difference in step number to grow. Although the unrestricted agent does converge to the optimal rewards in fewer episodes, it clearly takes more steps to do so because it has roughly 16 times the number of trainable parameters as the inductive bias agent has. The fact that we get comparable performance from the inductive bias agent using far fewer parameters justifies the relatively small fluctuations around the optimal value for each of the reward curves, which are expected to decrease as more agents are included in the average.
@@
VI A Quantum Motivation of the Inductive Bias and a Path Towards Quantization
In Section IV, we gave a qualitative, conceptual motivation of our inductive biases. There, we identified that the size of the power set $\mathcal{P}(V)$ can lead to exponential costs, because in general all of its elements might be valid excitation configurations for which we have to define $h$-values.
To get around this issue, we took a look at many-body physics, in which fundamental scattering processes typically only involve a handful of particles. This guided us to formulate a classical inductive bias in which complex transitions from observations to actions are given by successions of few-body transitions.
However, the analogy to quantum many-body physics can be worked out more formally, and this in particular suggests a natural quantization for analog quantum computers. Therefore, in Section VI.1 we will work out this formal analogy, while in Section VI.2 we will provide some basic considerations for an actual quantum implementation.
VI.1 Full Quantum Motivation
In this section, we will provide more formal details on the analogy to quantum many-body physics which inspired our inductive biases. For that purpose, we start by explaining how excitation configurations are represented in quantum physics. Specifically, each atomic clip in our hypergraph can be interpreted as a mode in quantum many-body physics and in the formalism called second quantization [32, 33], each excitation configuration is represented by a vector |$\ket{n_{1},\dots,n_{|V|}}$ in a Hilbert space, where $n_{j}$ is the number of excitations in atomic clip/mode $j$.
Time evolution over the time duration $\Delta t$ is described by an operator $U(\Delta t)=e^{-i\Delta t\cdot H}$, where $i$ is the complex unit, $H$ is an operator called the Hamiltonian, and $e^{\bullet}$ is the matrix exponential. This means after a duration $\Delta t$, a many-body system starting in a state |$\ket{n_{1},\dots,n_{|V|}}$ is afterwards described by the state $e^{i\Delta t\cdot H}\ket{n_{1},\dots,n_{|V|}}$.
Important for our inductive bias are the typical expressions for $H$. For this, we require the ladder operators $a_{j}$ and $a^{\dagger}_{j}$, with denoting the hermitian adjoint (i.e. $A^{\dagger}=\overline{A^{T}}$, with $\overline{\bullet}$ denoting complex conjugation) and $j\in V$. $a^{\dagger}_{j}$ adds one excitation to atomic clip $j$, i.e. $a^{\dagger}_{j}\ket{\ldots,n_{j},\ldots}\propto\ket{\ldots,n_{j}+1,\ldots}$ and is called a creation operator. Similarly, $a_{j}$ removes an excitation from an atomic clip $j$, i.e. $a_{j}\ket{\ldots,n_{j},\ldots}\propto\ket{\ldots,n_{j}-1,\ldots}$, and is called an annihilation operator. For the special case $n_{j}=0$, we have $a_{j}\ket{\ldots,n_{j},\ldots}=0$.
A typical Hamiltonian $H$ in second quantization is of the form
For small enough time intervals $\delta t$, the time evolution operator can be approximated as $e^{iH\delta t}\approx\mathbb{1}+iH\delta t$. We discard the identity operator $\mathbb{1}$ (in physical terms, we post-select on a non-trivial change occurring) because it means that no transition occurred at all. We discretize time in multiples of $\delta t$, and identify each step in the random walk with one application of $i\delta tH$, absorbing $i\delta t$ into the coefficients $h(j_{1},\ldots,j_{i},k_{1},\ldots,k_{o})$.
[…]
The last ingredient in our inductive bias is the observation that the operators $h(j_{1},\ldots,j_{i},k_{1},\ldots,k_{o})a^{\dagger}_{k_{1}}\cdots a^{\dagger}_{k_{o}}\cdot a_{j_{1}}\cdots a_{j_{i}}$ also act on excitation configurations |$\ket{n_{1},\ldots,n_{|V|}}$ that have excitations $n_{x}=1$ for $x\neq j_{1},\ldots,j_{i},k_{1},\ldots,k_{o}$, but leave $n_{x}$ unchanged. Similarly, we also apply $h$-values $h_{(i,o)}(\{c_{j_{1}},\ldots,c_{j_{i}}\},\{c_{k_{1}},\ldots,c_{k_{o}}\})$ to excitation configurations that have additional excitations in unrelated atomic clips $c_{x}$.
VI.2 First Steps Towards Quantum mePS